A projectile is launched from the ground at 30 degrees from te horizontal at 45 m/s. A target is placed 30 m horizontally from the starting point at 8 m above the ground. Is the projectile going to hit the target? if not by what vertical distance is it missing the target? Is the projectile moving upward or downward at that instance?
Answer #1

X = Vcos(theta)*t

=> 30 = 45*cos(30)*t

=> 30 = 45*0.866*t

=> t = 0.7698 sec

Y = ut + 1/2at^2 ( u = 45sin(30) , a = -9.8 m/s^2)

=> Y = 0.5*45 - 1/2*9.8*(0.7698)^2 = 19.59 m

=> it won't heat the target ...it will be above the target by 19.59-8 = 11.59 m